two bodies of mass 40 kg and 80 kg are at distance 0.15 m from each other foce is 1.0 mg - wt calculate G if g =10 m / s^2

Grade:11

1 Answers

jitender
6 years ago
F=1mg-wt=1×10^-6 NAnd r=0.15mM1=40 and M2=80F=GM1M2/r^2G=(Fr^2)/M1M2G=[10^-6 (0.15)^2]/40×80G=7.03×10^-12